# Black-box Membership Inference Attacks against Fine-tuned Diffusion Models

### v1

@@ 4.3 Competitors
In this scenario, the attacker does not possess a conditional prompt that can be directly fed into the target model. As illustrated in Figure 3, the attack scenario necessitates the use of an image captioning model to produce a caption for $I_{q}$. This caption is subsequently used as the input for $\mathcal{M}$. The process culminates in the computation of similarity between the query image $I_{q}$ and the image generated by $\mathcal{M}$.
@@
is similar to the second scenario.
Figure 4: Relationship between epoch progression and attack success rates in  Attack-IAttack-IIAttack-III, and Attack-IV, indicating increasing memorization within image generation models over fine-tuning epochs.
5 Experiments Evaluation
5.1 Comparison with Baselines
TABLE II: Accuracy comparison between the attacks by Zhang et al. [63] and Matsumoto et al. [34] versus our methods: Attack-IAttack-IIAttack-III, and Attack-IV, illustrating the distinct success rates.
@@ 5.1 Comparison with Baselines
Results are shown in Table II. We ensure consistency in simulating real-world scenarios, wherein the number of images that a malicious publisher can sample from the target generator is limited. Under the constraint of limited sample size, we observe that the accuracy of both baseline attacks nearly equates to random guessing. We conjecture that this is due to their reliance on a large number of synthesis images for decision-making. Specifically, Zhang et al. [63]’s approach requires learning the distributional differences between generated image samples and non-member samples using ResNet18, based on a substantial volume of images sampled from the target model, and subsequently applying this knowledge to assess the input query data. However, such an attack premise falters in realistic scenarios where a malicious model publisher restricts the number of images a user can obtain from the model, preventing attackers from sampling a large volume of images to conduct the attack. Under such constraints, the effectiveness of attacks by Zhang et al. [63] and others is inevitably compromised, as the insufficient sample size hampers the ability to accurately discern the differences between the two data distributions. Similarly, the approach by Matsumoto et al. [34] encounters a hurdle; in scenarios of limited generative sample availability, it becomes challenging to find a suitable reconstruction counterpart and calculate its distance from the original data record. Consequently, these methods fail to achieve high attack success rates under sample-restricted conditions. In contrast, the four attack methods we propose still attain a high success rate despite the limited number of generative samples. This is attributed to our attacks being based on the similarity scores as proposed in Section 3.1, which, while influenced by the quality of the model’s generated images, is not hindered by the quantity of these images.
@@
5.2 Different Image Encoder
@@
As depicted in Figure 2, our five image feature extractors excel across four different attack scenarios within the classifier-based attack domain. Each maintains an attack accuracy exceeding $70\%$, underscoring the robustness of our attack framework across different feature extractors. Notably, the implementation of DeiT [59] as the feature extraction model yielded a marginally higher and more consistent success rate compared to the other image encoders. Therefore, we selected DeiT as the default image encoder for future experimental investigations.
A more comprehensive comparison including threshold-based and distribution-based of these five image encoders is presented in Appendix F.
5.3 Different Distance Metrics
@@
5.4 Impact of Fine-tuning Steps
We then investigated the influence of the number of fine-tuning steps on the success rate of attacks. Evaluations were conducted at intervals of $100$ epochs, ranging from $100$ to $500$ epochs, to measure the attack success rate. As the model’s memorization of the training data can be equated to overfitting effects, it is anticipated that with an increased number of fine-tuning steps, the model increasingly exhibits a tendency towards overfitting and enhanced memorization of the training samples. Consequently, when querying the model with member set samples compared to non-member samples, a more distinct similarity discrepancy should be observed.
@@ 5.4 Impact of Fine-tuning Steps
We have also included the results for threshold-based and distribution-based attacks under these four scenarios in the Appendix D for reference.
5.5 Impact of Number of Inference Step
@@
TABLE III: Alignment with DDIM [54] denoting ‘S’ as inference steps; experimentation under Attack-III scenario measuring FID at varying inference step counts.
@@
30 $0.75$ $0.8225$ $0.30$ $0.76$ $0.8816$ $0.50$ $0.865$ $0.93$ $0.54$ $8.77$
50 $0.765$ $0.8146$ $0.25$ $0.77$ $0.8920$ $0.37$ $0.85$ $0.93$ $0.58$ $7.835$
100 $0.74$ $0.8172$ $0.26$ $0.745$ $0.8818$ $0.40$ $0.855$ $0.94$ $0.61$ $7.527$
200 $0.745$ $0.8125$ $0.39$ $0.74$ $0.8869$ $0.49$ $0.87$ $0.94$ $0.58$ $7.472$
@@
As illustrated in Table III, the variations in attack accuracy are not immediately pronounced. However, upon a broader examination, it becomes evident that as the number of S (inference steps) increases, there is a gradual uptrend in the success rate of attacks. Notably, attacks based on classifiers yield the highest accuracy. To delve deeper into the reason why an increased number of inference steps doesn’t lead to a substantial boost in attack success rate, we present samples generated at different inference steps in Figure 6. It becomes apparent that as the number of inference steps rises, only certain localized features of the generated images are affected. The overall style remains largely undisturbed, with no significant discrepancies observed. This observation potentially explains why altering the inference steps doesn’t drastically impact the attack success rate.
The experimental results obtained from the additional two datasets are presented in Appendix E.
5.6 Impact of Different Size of Auxiliary Dataset
From our observations across white-box [6, 34, 20], gray-box [12, 25, 20], and black-box attacks [34], the accuracy of these attacks is significantly influenced by the size of training set. As the training set of the target model, $\mathcal{M}^{t}$, encompasses more samples, its “memorization” capability for individual samples diminishes. This is attributed to the fact that an increase in training data can decelerate the model’s convergence rate, impacting its ability to fit all the training sets accurately. As a result, many attacks do not demonstrate effective performance as the dataset size expands. In this work, we investigate how increasing the size of the dataset used by the target model affects the success rate of our black-box attack. Given that our work is predicated on leveraging pre-trained models for downstream tasks, where the downstream datasets usually do not contain a vast number of samples, we have established our training dataset sizes at $100$, $200$, $500$, and $1000$. Using the CelebA dataset, we aim to assess the variations in the performance of the three attack types when the attacker is privy to four distinct values of knowledge.
5.7 Impact of the Selection of Shadow Models
To enhance the universality and applicability of our attack methodology in real-world scenarios, we propose to further relax the assumptions pertaining to the attack environment. In our prior experiments, all results were predicated on the use of shadow models mirroring the target model’s structural framework to generate training data for the attack inference model. However, in practical settings, malicious model publishers may withhold any specific details about the model, offering only a user interface. Under such circumstances, it is not advisable to confine ourselves to a specific type of shadow model. Instead, a more effective approach would be to leverage the memorization properties of image generators when creating training data for the attack, thus diversifying and strengthening the attack strategy.
Therefore, we will employ a conditional image generator, Kandinsky [42], which has a different architectural design from the Stable Diffusion [44], as our shadow model. This model will be fine-tuned using the same auxiliary dataset mentioned in the Section 5 and the results will be displayed in Table IV.
TABLE IV: Use of Kandinsky [42] as shadow model and Stable Diffusion [44] as target model in conducting attacks, demonstrating the maintained efficacy of all four attack methodologies.
@@ 5.7 Impact of the Selection of Shadow Models
In Table IV, we evaluate attackers with different knowledge across three datasets, employing an MLP model as the attack inference model. The notation ‘*-*-S’ indicates attacks conducted using a shadow model with the same architecture as the target model. Conversely, ‘*-*-A’ denotes scenarios where the target model is anonymous to the attacker, hence the shadow model and the target model are architecturally dissimilar. The experimental data indicate that altering the shadow model has only a minimal effect on the success rate of the attacks, with all attacks still capable of achieving a relatively high level of success. This further substantiates the robustness and generalizability of our attack framework.
5.8 Impact of Eliminating Fine-Tuning in Captioning Models
In our work, within the attack environments designed for Attack-II and Attack-IV, the attacker does not have full access to the query point $x$, but only a query image $I_{q}$. In Section 5, for these two attack scenarios, we initially use auxiliary data to fine-tune the image captioning model before generating matching prompt information based on the query image. However, this approach significantly increases the time cost of the attack. Therefore, we use an image captioning model that has not been fine-tuned to generate image descriptions. We then carry out the attack based on these generated descriptions.
TABLE V: Impact of not fine-tuning the captioning model on the success rates of Attack-II and Attack-IV across various datasets.
@@ 5.8 Impact of Eliminating Fine-Tuning in Captioning Models
From Table V, it is evident that without fine-tuning the captioning model, there is a varying degree of reduction in the success rates of attacks across different datasets. Notably, when using CelebA-Dialog as the test set, the success rate of the attack drops by nearly $30\%$, leading to a marked inconsistency in the attack outcomes. Unlike changing the types of shadow models, a captioning model without tuning more conspicuously diminishes the effectiveness of the attacks. We posit the reason is the image captioning model may have introduced biases in the generated $T_{q}$, adversely affecting the quality of the resultant images.
Takeaways:  We compared the four attack scenarios we proposed with existing black-box attacks and found that our accuracy significantly surpasses the established baselines. To thoroughly evaluate the accuracy and stability of our attacks, we conducted tests employing various image encoders, distance metrics, fine-tuning steps, and inference procedures, as well as different sizes of auxiliary datasets. Additionally, we experimented with changing the types of shadow models and testing without fine-tuning the image caption model. Our findings reveal a strong correlation between the attacks’ success rate and the generated images’ quality. Higher quality images lead to increased attack success rates, which aligns with the theory of similarity scores mentioned in Section 3.1.
6 Defense
@@
Figure 8: Results of attacks with DP-SGD: Despite the implementation of defenses, the attacks still retain partial effectiveness.
We employed a classifier as the attack inference model to test against the MS COCO dataset. The results indicate that the success rate of the attack significantly decreases after implementing DP-SGD [1] as a defensive measure. Figure 8 illustrates that the accuracy of Attack-III and Attack-IV is significantly impacted, dropping to around $60\%$, whereas Attack-I and Attack-II still maintain a success rate close to $70\%$.
We hypothesize that our attack relies on the model’s memorization of training data, which enables the use of similarity scores as an indicator for assessment. However, the application of DP-SGD reduces the model’s memorization capacity for training data, leading to decreased similarity scores in member set samples and, resulting in a lower attack success rate. The higher attack accuracy of Attack-I and Attack-II can be attributed to the overlap between the member sets of the shadow models and the training set of the target models.
7 Related Work
We review further related works on membership inference attacks (MIAs) and extraction attacks against diffusion models.
7.1 White-Box MIA
@@ 8 Conclusion
However, our attacks still face certain limitations. As discussed in Section 5.8, both Attack-II and Attack-IV critically rely on a captioning model that has been fine-tuned using an auxiliary dataset. We hope future work can effectively address this challenge.
@@ Appendix A More Details for Diffusion Models
Given that the original work [18] considers the forward process as a Markov chain and employs the reparameterization trick, it is possible to directly derive the noisy image $x_{t}$ at step $t$ from the original image $x_{0}$.
Given the definition $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, the training objective for diffusion models is to obtain a denoising network capable of sampling $x_{t-1}$ from $x_{t}$ according to the distribution $\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\sigma_{t}^{2}\text{I})$. When conditioned on $x_{0}$ and $x_{t}$ from $q(x_{t-1}|x_{t},x_{0})$, the ground-truth distribution of $x_{t-1}$ is given by $\mathcal{N}(x_{t-1};\tilde{\mu}_{t}(x_{t},x_{0}),\sigma_{t}^{2}\text{I})$. Given that the variance is fixed as a hyperparameter, the focus is on calculating the difference between $\mu_{\theta}(x_{t},t)$ and $\tilde{\mu}_{t}(x_{t},x_{0})$. Applying Bayes’ rule to the ground-truth distribution

### v4

@@ IV-C Competitors
In this scenario, the attacker does not possess a conditional prompt that can be directly fed into the target model. As illustrated in Figure 2, the attack scenario necessitates the use of an image captioning model to produce a caption for $I_{q}$. This caption is subsequently used as the input for $\mathcal{G}$. The process culminates in the computation of similarity between the query image $I_{q}$ and the image generated by $\mathcal{G}$.
@@
is similar to the second scenario (there is no overlap between the attacker’s auxiliary dataset and the target member set). This attack represents the hardest situation, and we think it will get the lowest accuracy.
V Experiments Evaluation
V-A Comparison with Baselines
TABLE III: Accuracy comparison between the attacks by Zhang et al. [63] and Matsumoto et al. [34] (Applying GAN-Leaks against the diffusion model) versus our methods: Attack-IAttack-IIAttack-III, and Attack-IV, illustrating the distinct success rates.
@@ V-A Comparison with Baselines
Results are shown in Table III. We ensure consistency in simulating real-world scenarios, wherein the number of images that a malicious publisher can sample from the target generator is limited. Under the constraint of limited sample size, we observe that the accuracy of both baseline attacks nearly equates to random guessing. We conjecture that this is due to their reliance on a large number of synthesis images for decision-making. Specifically, Zhang et al. [63]’s approach requires learning the distributional differences between generated image samples and non-member samples using ResNet18, based on a substantial volume of images sampled from the target model, and subsequently applying this knowledge to assess the input query data. However, such an attack premise falters in realistic scenarios where a malicious model publisher restricts the number of images a user can obtain from the model, preventing attackers from sampling a large volume of images to conduct the attack. Under such constraints, the effectiveness of attacks by Zhang et al. [63] and others is inevitably compromised, as the insufficient sample size hampers the ability to accurately discern the differences between the two data distributions. Similarly, the approach by Matsumoto et al. [34] encounters a hurdle; in scenarios of limited generative sample availability, it becomes challenging to find a suitable reconstruction counterpart and calculate its distance from the original data record. Consequently, these methods fail to achieve high attack success rates under sample-restricted conditions. In contrast, the four attacks we propose still attain a high success rate despite the limited number of generative samples. This is attributed to our attacks being based on the similarity scores as proposed in Section III-A, which, while influenced by the quality of the model’s generated images, is not hindered by the quantity of these images.
@@
Figure 6: Relationship between epoch progression and AUC score in Attack-IAttack-IIAttack-III, and Attack-IV, indicating increasing memorization within image generation models over fine-tuning epochs.
V-B Different Image Encoder
@@
As depicted in Figure 3, our five image feature extractors excel across four different attack scenarios within the classifier-based attack domain. Each maintains an AUC score exceeding $0.7$, underscoring the robustness of our attack framework across different feature extractors. Notably, the implementation of DeiT [59] as the feature extraction model yielded a marginally higher and more consistent success rate compared to the other image encoders. Therefore, we selected DeiT as the default image encoder for future experimental investigations.
A more comprehensive comparison including threshold-based and distribution-based of these five image encoders is presented in Appendix E.
V-C Different Distance Metrics
@@
V-D Impact of Fine-tuning Steps
We then investigated the influence of the number of fine-tuning steps on the success rate of attacks. Evaluations were conducted at intervals of $100$ epochs, ranging from $100$ to $500$ epochs, to measure the attack success rate. The default image encoder and distance metrics are Deit and Cosine similarity; all fine-tuning settings are aligned with Table II. As the model’s memorization of the training data can be equated to overfitting effects, it is anticipated that with an increased number of fine-tuning steps, the model increasingly exhibits a tendency towards overfitting and enhanced memorization of the training samples. Consequently, when querying the model with member set samples compared to non-member samples, a more distinct similarity discrepancy should be observed.
@@ V-D Impact of Fine-tuning Steps
We have also included the results for threshold-based and distribution-based attacks under these four scenarios in the Appendix F for reference.
V-E Impact of Number of Inference Step
@@
TABLE IV: Alignment with DDIM [54] denoting ‘S’ as inference steps; experimentation under Attack-III scenario measuring FID at varying inference step counts.
@@
$30$ $0.75$ $0.8225$ $0.30$ $0.76$ $0.8816$ $0.50$ $0.865$ $0.93$ $0.54$ $8.77$
$50$ $0.765$ $0.8146$ $0.25$ $0.77$ $0.8920$ $0.37$ $0.85$ $0.93$ $0.58$ $7.835$
$100$ $0.74$ $0.8172$ $0.26$ $0.745$ $0.8818$ $0.40$ $0.855$ $0.94$ $0.61$ $7.527$
$200$ $0.745$ $0.8125$ $0.39$ $0.74$ $0.8869$ $0.49$ $0.87$ $0.94$ $0.58$ $7.472$
As illustrated in Table IV, the variations in attack accuracy are not immediately pronounced. However, upon a broader examination, it becomes evident that as the number of S (inference steps) increases, there is a gradual uptrend in the success rate of attacks. Notably, attacks based on classifiers yield the highest accuracy. To delve deeper into the reason why an increased number of inference steps doesn’t lead to a substantial boost in attack success rate, we present samples generated at different inference steps in Figure 7. It becomes apparent that as the number of inference steps rises, only certain localized features of the generated images are affected. The overall style remains largely undisturbed, with no significant discrepancies observed. This observation potentially explains why altering the inference steps doesn’t drastically impact the attack success rate.
The experimental results obtained from the additional two datasets are presented in Appendix G.
V-F Impact of Different Size of Auxiliary Dataset
From our observations across white-box [6, 34, 20], gray-box [12, 25, 20], and black-box attacks [34], the accuracy of these attacks is significantly influenced by the size of training set. As the training set of the target model, $\mathcal{G}^{t}$, encompasses more samples, its “memorization” capability for individual samples diminishes. This is attributed to the fact that an increase in training data can decelerate the model’s convergence rate, impacting its ability to fit all the training sets accurately. As a result, many attacks do not demonstrate effective performance as the dataset size expands. In this work, we investigate how increasing the size of the dataset used by the target model affects the success rate of our black-box attack. Given that our work is predicated on leveraging pre-trained models for downstream tasks, where the downstream datasets usually do not contain a vast number of samples, we have established our training dataset sizes at $100$, $200$, $500$, and $1000$. Using the CelebA dataset, we aim to assess the variations in the performance of the three attack types when the attacker is privy to four distinct values of knowledge.
@@ V-F Impact of Different Size of Auxiliary Dataset
V-G Impact of the Selection of Shadow Models
To enhance the universality and applicability of our attack methodology in real-world scenarios, we propose to further relax the assumptions pretraining to the attack environment. In our prior experiments, all results were predicated on the use of shadow models mirroring the target model’s structural framework to generate training data for the attack inference model. However, in practical settings, malicious model publishers may withhold any specific details about the model, offering only a user interface. Under such circumstances, it is not advisable to confine ourselves to a specific type of shadow model. Instead, a more effective approach would be to leverage the memorization properties of image generators when creating training data for the attack, thus diversifying and strengthening the attack strategy.
Therefore, we will employ a conditional image generator, Kandinsky [42], which has a different architectural design from the Stable Diffusion [44], as our shadow model. This model will be fine-tuned using the same auxiliary dataset mentioned in the Section IV-A, and the results will be displayed in Table V.
TABLE V: Use of Kandinsky [42] as shadow model and Stable Diffusion [44] as target model in conducting attacks, demonstrating the maintained efficacy of all four attack methodologies.
@@ V-G Impact of the Selection of Shadow Models
In Table V, we evaluate attackers with different knowledge across three datasets, employing a classifier as the attack inference model. The notation ‘*-*-S’ indicates attacks conducted using a shadow model with the same architecture as the target model. Conversely, ‘*-*-A’ denotes scenarios where the target model is anonymous to the attacker, hence the shadow model and the target model are architecturally dissimilar. The experimental data indicate that altering the shadow model has only a minimal effect on the success rate of the attacks, with all attacks still capable of achieving a relatively high level of success. This further substantiates the robustness and generalizability of our attack framework.
V-H Impact of Eliminating Fine-Tuning in Captioning Models
In our work, within the attack environments designed for Attack-II and Attack-IV, the attacker does not have full access to the query point $x$, but only a query image $I_{q}$. In previous sections, for these two attack scenarios, we initially used auxiliary data to fine-tune the image captioning model before generating matching prompt information based on the query image. However, this approach significantly increases the time cost of the attack. Therefore, we use an image captioning model that has not been fine-tuned to generate image descriptions. We then carry out the attack based on these generated descriptions.
TABLE VI: Impact of not fine-tuning the captioning model on the success rates of Attack-II and Attack-IV across various datasets.
@@ V-H Impact of Eliminating Fine-Tuning in Captioning Models
From Table VI, it is evident that without fine-tuning the captioning model, there is a varying degree of reduction in the success rates of attacks across different datasets. Notably, when using CelebA-Dialog as the test set, the success rate of the attack drops by nearly $30\%$, leading to a marked inconsistency in the attack outcomes. Unlike changing the types of shadow models, a captioning model without tuning more conspicuously diminishes the effectiveness of the attacks. We posit the reason is the image captioning model may have introduced biases in the generated $T_{q}$, adversely affecting the quality of the resultant images.
TABLE VII: Explore the attack accuracy under DP-SGD defense—our four attack methods’ accuracy declines. Experiments include two different sizes of fine-tuning datasets and three $\epsilon$ values. ‘Vanilla’ means without the DP-SGD version.
Attack-I Attack-II Attack-III Attack-IV
ASR AUC T@1%F ASR AUC T@1%F ASR AUC T@1%F ASR AUC T@1%F
$100$ $\epsilon=1$ $0.581$ $0.646$ $0.01$ $0.532$ $0.654$ $0.01$ $0.495$ $0.498$ $0.00$ $0.522$ $0.524$ $0.00$
$\epsilon=4$ $0.592$ $0.651$ $0.01$ $0.575$ $0.647$ $0.01$ $0.515$ $0.514$ $0.01$ $0.535$ $0.534$ $0.01$
$\epsilon=10$ $0.595$ $0.641$ $0.02$ $0.560$ $0.644$ $0.02$ $0.56$ $0.522$ $0.01$ $0.545$ $0.522$ $0.01$
Vanilla $0.843$ $0.911$ $0.58$ $0.845$ $0.909$ $0.51$ $0.831$ $0.893$ $0.38$ $0.665$ $0.713$ $0.12$
$200$ $\epsilon=1$ $0.593$ $0.632$ $0.01$ $0.628$ $0.676$ $0.01$ $0.493$ $0.502$ $0.00$ $0.548$ $0.524$ $0.01$
$\epsilon=4$ $0.601$ $0.652$ $0.01$ $0.618$ $0.670$ $0.01$ $0.523$ $0.516$ $0.01$ $0.515$ $0.506$ $0.01$
$\epsilon=10$ $0.585$ $0.632$ $0.03$ $0.643$ $0.655$ $0.02$ $0.535$ $0.504$ $0.01$ $0.542$ $0.541$ $0.02$
Vanilla $0.767$ $0.863$ $0.30$ $0.730$ $0.812$ $0.11$ $0.695$ $0.728$ $0.09$ $0.673$ $0.700$ $0.05$
Takeaways:  We compared the four attack scenarios we proposed with existing black-box attacks and found that our accuracy significantly surpasses the established baselines. To thoroughly evaluate the accuracy and stability of our attacks, we conducted tests employing various image encoders, distance metrics, fine-tuning steps, and inference procedures, as well as different sizes of auxiliary datasets. Additionally, we experimented with changing the types of shadow models and testing without fine-tuning the image caption model to test our attacks’ generalization and robustness. Our findings reveal a strong correlation between the attacks’ success rate and the generated images’ quality. Higher quality images lead to increased attack success rates, which aligns with the theory of similarity scores mentioned in Section III-A.
VI Defense
@@
In this part, we test our four attacks and employ a classifier as the inference model against the MS COCO dataset. Due to the limit of time and computing resources, we only selected two different sizes of datasets, $100$ and $200$. For the DP-SGD mechanism, we set clipping norm $C=1$, $\delta=1\times 10^{-3}$, sampling rate $q=4/$(dataset size), epoch number is $500$, and target privacy budget $\epsilon\in\{1,4,10\}$ (different $\epsilon$ gives different noise multiplier $\sigma$).
We show the experimental results in Table VII. It illustrates that the attack success rate of four attacks significantly decreases after implementing DP-SGD [1] as the defensive method. When we use $100$ samples to fine-tune the model and set $\epsilon=1$, both Attack-III and Attack-IV have been greatly impacting, dropping to around $50\%$ (random guess). For Attack-I and Attack-II, both attacks allow attackers to be able to access the target model’s data sample partially. Therefore, it can still get over $55\%$ attack accuracy and seems robust to the defense. We also measure attack success rate at low FPRs, according to Carlini et al. [5]. From Table VII, we noticed the TPR value when FPR=$1\%$ has dropped from $0.58$ and $0.51$ to $0.01$ and $0.01$. This phoneme shows that Attack-I and Attack-II also lose their functionality in these defense settings.
When we change the $\epsilon$ value from $1$ to $4$ and $10$, the attack accuracy increases but still cannot show their effectiveness. We hypothesize that our attack relies on the model’s memorization of training data, which enables the use of similarity scores as an indicator for assessment. However, the application of DP-SGD reduces the model’s memorization capacity for training data, leading to decreased similarity scores in member set samples and, resulting in a lower attack success rate. The experiment outcome has aligned with other works [6, 20, 12].
VII Related Work
We review further related works on membership inference attacks and extraction attacks against diffusion models.
VII-A White-Box MIA
@@ VIII Conclusion
However, our attacks still face certain limitations. As discussed in Section V-H, both Attack-II and Attack-IV critically rely on a captioning model that has been fine-tuned using an auxiliary dataset. We hope future work can effectively address this challenge.
@@ Appendix A More Details for Diffusion Models
Given that the original work [18] considers the forward process as a Markov chain and employs the reparameterization trick, it is possible to directly derive the noisy image $x_{t}$ at step $t$ from the original image $x_{0}$. Based on the definition $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$, the training objective for diffusion models is to obtain a denoising network capable of sampling $x_{t-1}$ from $x_{t}$ according to the distribution $\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\sigma_{t}^{2}\text{I})$. When conditioned on $x_{0}$ and $x_{t}$ from $q(x_{t-1}|x_{t},x_{0})$, the ground-truth distribution of $x_{t-1}$ is given by $\mathcal{N}(x_{t-1};\tilde{\mu}_{t}(x_{t},x_{0}),\sigma_{t}^{2}\text{I})$. Given that variance is fixed as a hyperparameter, the focus is on calculating the difference between $\mu_{\theta}(x_{t},t)$ and $\tilde{\mu}_{t}(x_{t},x_{0})$. Applying Bayes’ rule to the ground-truth distribution